Can you please help me with the following 2 questions.
Thanks
1.
A = P + PRT
SOLVE FOR T WHEN A = 70, P = 50 AND R = .1
2. FIND THE 3 CONSECUTIVE INTEGERS SUCH THAT SMALLER OF THE THREE (3) IS 61 MORE THAN THE LARGEST.

Answer by carl(7)

i can help with the first part....
ok.. first lets rite the question.... A = P + PRT and then let's plug in the givens...
70 = 50 + (50 x 1.2 x T) now 50 x 1.2 = 60T so lets rewrite the problem...
70 = 50 + 60T and now we are gonna subtract 50 from both sides and we get
20 = 60T ..... now divide out the 60 on both sides and we get 20/60 and 60/60T and we get
1/3 = T..... now just to be sure lets plug it in and see if it works...
A = 50 + (50 x 1.2 x .33)
A = 70
Good to go???
Carl